# Motion2VecSets: 4D Latent Vector Set Diffusion for Non-rigid Shape Reconstruction and Tracking

### 12 Jan 2024 v1

@@ 4.2.1 Monocular Depth Sequences
To simulate sparse and partial real-world scans, we generated monocular depth sequences by rendering from a fixed camera angle. The size of the point cloud input and the frame length are the same as Sec. 4.1.
The qualitative and quantitative comparisons are presented in Figure 6 and Table 2, respectively. As seen, our method consistently outperforms all state-of-the-art methods in all metrics and produces more complete surface geometries with more plausible motion tracking. This demonstrates the effectiveness of the motion priors learned by our proposed 4D latent set diffusion in addressing ambiguous data like partial observations.
@@ 4.2.2 Partial Scan Sequences
To assess the robustness of our method to extremely ambiguous data, we set up a challenging experiment on the D-FAUST [4] dataset. This involved reconstructing whole body motions based on partial point clouds of the upper bodies. This setup creates a highly ambiguous scenario, as the same upper body motion can correspond to many different lower-body. We adopt the same configuration as Sec. 4.1, with a frame length () and input point cloud size (). OFlow [36], LPDC [52], and CaDex [26] encounter challenges in reconstructing the complete shape, often producing distorted shapes such as broken feet. In contrast, our method excels in reconstructing more complete geometries while achieving temporally coherent tracking. Additionally, our approach present a diverse range of plausible full-body reconstructions that align with the given upper-body scans. The superior performance of our model is primarily attributed to the 4D latent vector set diffusion. Our diffusion-based method is more capable of tackling the ‘one-to-many’ complexities from extremely partial data.
@@ 4.3 Ablation Study
We conducted detailed ablation studies to validate the effectiveness of each design (See Table 3).
@@
Instead of using global latent codes to represent shapes and time-varying deformations, our approach employs 4D latent vector sets. As indicated by Table 3 (with ), our method significantly outperforms the variant using global latent codes, achieving a 15.3 improvement in IoU for unseen motion reconstruction. This highlights the superior ability of 4D latent sets to accurately capture 4D motions. The advantage becomes even more apparent with a 16.7 % improvement in unseen identities, demonstrating enhanced generalization ability to objects with different structures.
@@
For the synchronized deformation latent set diffusion, we have integrated the time self-attention layer in our interleaved spatio-temporal attention mechanism. We attempted to remove the layer. However, the results exhibited a decrease in all metrics, emphasizing the effectiveness of the time self-attention layer in maintaining temporal coherence.
@@
To determine the optimal configuration for learning shape and deformation priors within time-varying deformable surfaces, we varied the channel numbers (8, 16, 32) of the shape and deformation latent sets. Experimental results indicated that using 32 channels for 4D latent set diffusion is more suitable, obtaining more favorable results.
@@ 5 Conclusion
We present Motion2VecSets, a 4D diffusion model for dynamic surface reconstruction from point cloud sequences. Our method explicitly models the shape and motion distribution of non-rigid objects through an iterative denoising process of compressed latent vector sets, generating plausible and diverse outputs. The learned shape and motion diffusion priors can effectively deal with ambiguous observations, including sparse, noisy, and partial data. Compared to encoding shape and deformation with a global latent, our novel 4D latent vector set representation enables more accurate non-linear motion capture and improves the generalizability to unseen identities and motions. The designed interleaved space and time attention block for synchronized deformation vector sets diffusion enforces temporal-coherent object tracking while reducing computational overhead. Extensive experiments on reconstructing sparse, partial, and even half bodies on the D-FAUST [4] and the DT4D-A [29] demonstrate the superiority of our approaches for various imperfect observations. We believe our Motion2VecSets can be future extended with multi-modalities such as text or speech-driven 4D generation.
@@ Acknowledgement.
In this supplementary material, we complement our main paper with additional details. It begins with an introduction to the notations used in our paper in Sec. A. The following section, Sec. B, discusses our network architectures, highlighting key design choices and their impact. This is complemented by Sec. C, which provides essential implementation details. We then show more comparisons between our approach with established methods like OFlow [36], LPDC [52], and CaDeX [26] in Sec. D, offering insights into our model’s strengths and improvements. Finally, in Sec. E, we demonstrate the real-world applicability of our model, underscoring its practical effectiveness.
@@ Appendix A Notations
In this paper, we use the following notations, as summarized in Tab. 4. The symbol represents the sequence length, with the superscript indicating the time step of a frame in the sequence. Our network’s input is a sparse or partial point cloud, denoted as , where each frame comprises points, serving as conditions in reference time. The sets represent shape codes, deformation codes, and conditional codes, respectively, with each -th code denoted by the corresponding lowercase letter. The mesh is symbolized by , where and correspond to the vertices and faces set of the mesh. Furthermore, refers to surface points sampled from ground truth meshes for learning shape, we dedicate for representing the sampled surface points for learning deformation. Moreover, represents the set of query points, where refers to a point inside the point set and is the output by feeding the query point into the deformation network.
Appendix B Network Archtectures
B.1 Shape Diffusion
@@
Points sample from meshes as input for shape autoencoder | |
Points sample from source and target meshes as input for deformation autoencoder | |
@@
In the Shape Autoencoder Network, points , surface and near-surface points sampled from the first frame’s mesh. To optimize computational efficiency, we adopt the furthest point sampling (FPS) technique. This method pinpoints crucial points within a point cloud, thereby thinning its density.:
@@
Nosised shape codes are sent to the denoising neural network. The denoiser consists of two blocks: The space-attention block facilitates positional information exchange among codes in different positions, while the condition-attention block injects information from sparse or partial points (conditional input). After repeating this process, a denoised latent set, is obtained.
@@ B.2 Synchronized Deformation Diffusion
As shown in Fig. 9, the deformation encoding part is also split into two parts: Deformation Autoencoder Network and Synchronized Deformation Vector Set Diffusion Network.
@@
We employ the cross attention operation throughout our method, where denotes the query, and denotes the key-value pair. Similarly, we use KL-divergence to retain high-level features to facilitate the learning of the diffusion model. Here we get the deformation latent set :
In inference time, we take the surface points of the source mesh as query points. The interaction between the predicted deformation codes sampled from the Shape Vector Set Diffusion Network and the positional embedding of query points through cross-attention yields the approximated latent features. Subsequently, a linear layer derives the predicted positions of the target point cloud.
@@
In this stage, the sparse noisy input point clouds pairs green blocks in Fig. 9 are processed with a conditional encoder. The conditional encoder has the same structure with the deformation encoder, through which we get the conditional latent set . In the meanwhile, the shape latent set from the deformation encoder is added with Gaussian noise and sent to the denoiser of the diffusion model. As illustrated in the figure, after the space self-attention block, the conditional latent set is injected into the cross-attention block, after the cross-attention along the temporal domain, we get the denoised deformation codes. This process is repeated for 18 times by default.
@@ C.1 Dataset
For shape encoding, our process begins with the application of the Butterfly Subdivision method [64], an interpolating subdivision technique widely used in computer graphics for generating smooth surfaces from polygonal meshes. Following the methodology outlined in [33], we center and rescale all meshes to ensure that the bounding box of each mesh is centered at the origin and the longest edge is normalized to a length of 1. We then uniformly sample 200k points within this normalized cube and compute their occupancy values to determine whether they lie inside or outside the mesh, as detailed in [33]. To enhance our model’s understanding of surface properties, we introduce Gaussian noise at two different levels to the mesh surface points. This process generates 200k near-surface points, whose occupancy values are also computed. These points, combined with an additional 200k points sampled directly from the mesh surface, constitute the input to our network, providing a comprehensive set of data points for both uniform space and near-surface regions.
@@
Within the context of the partial challenge configuration, a camera was positioned with a fixed viewing angle of degrees directed towards the human or animal subject within the scene. This positioning was undertaken to capture a partial depth observation of the subject, as illustrated in Figure 10. Additionally, in the more restricted partial setting, only half of the human body was observed. This was accomplished by intentionally selecting vertex indices corresponding to the upper body of the SMPL model [31].
@@ Appendix D Additional Results
D.1 4D Shape Reconstruction
In evaluating our model’s performance in 4D shape reconstruction, we adopted a comprehensive set of metrics: Intersection over Union (IoU), chamfer distance, and correspondence distance. These metrics were chosen for their relevance in accurately quantifying shape reconstruction quality. IoU measures the overlap between predicted and ground truth shapes, Chamfer distance quantifies the average closest point distance, and correspondence distance evaluates the accuracy of point-wise correspondences. These metrics align with the standards set in recent studies, such as those by LPDC [52] and OFlow [36].
We present the average metrics across all 17 frames for both D-FAUST [4] and DT4D-A [29] datasets. The quantitative results are shown in Tab. 6, Tab. 5, Tab. 8, and Tab. 7. Our analysis reveals that our model demonstrates superior generalization capabilities in scenarios involving both unseen motions and unseen individuals. Specifically, it outperforms existing approaches, including OFlow [36], LPDC [52], and CaDeX [26], across all evaluated metrics for all time frames. This advancement underscores the efficacy of our approach in handling the complexities of 4D shape reconstruction. Furthermore, qualitative results in Fig. 12 and Fig. 13 show a selection of 8 frames, chosen to represent a diverse range of motions and shapes, from the total 17 to illustrate our model’s performance. In each figure, the upper part displays the results for unseen motion, while the lower part corresponds to unseen individuals. We utilize a chamfer distance error map for visualization, where blue indicates lower error and red signifies higher error. The color-coded error map, computed based on the distance between predicted and ground truth points, provides an intuitive understanding of the model’s accuracy in different scenarios. Our model not only has an overall smaller error on both datasets, but also captures motions more accurately.
D.2 4D Shape Completion
In the more challenging task of 4D shape completion from partial point clouds, our model shows noticable improvements over existing state-of-the-art methods. The quantitative results, as illustrated in Fig. 14 and Fig. 15, demonstrate substantial performance enhancements, with a lower error rate in scenarios involving unseen motions and individuals. This underscores the robustness of our approach.
Our model’s capabilities are especially evident in the D-FAUST dataset [4], achieving more accurate reconstructions of complex human body parts, like hands and feet. These parts, typically difficult to reconstruct, are rendered with higher fidelity.
Furthermore, the DT4D-A dataset [29], known for its varied topologies, shows our model’s strengths. In this dataset, our model not only demonstrates lower reconstruction errors for animal shapes but also captures their motion with greater accuracy. This performance, across a diverse range of topologies, highlights the adaptability and effectiveness of our model in complex 4D shape reconstruction tasks.
Appendix E Real-world Data Test
In this section, we validate our model using data from the real-world BEHAVE dataset [2], employing four Kinect RGB-D cameras to capture RGB color and depth frames. In our case, we utilize a single view from a fixed camera to align with our previous partial scan setting. Similarly, the depth map is back-projected into a 3D point cloud, serving as a partial input for our model. In Fig. 11, we present a visual representation of our reconstruction process with the corresponding input and reference RGB frame.
The results demonstrate the robustness of our model in scenarios characterized by markedly incomplete scans, such as instances where limbs, like the leg or arm, are obscured by structures such as a grasped object in the hand. This resilience stems from the inherent capabilities of the diffusion model, empowering our model to infer potential structures even in the presence of significant occlusions. This effectiveness is attributed to the learned latent distribution, highlighting the model’s capacity to generalize and extrapolate beyond the limitations imposed by incomplete or obstructed input scans.
Notably, our method exhibits remarkable reliability even without the need for model fine-tuning on the real dataset. This stability is particularly evident when addressing inherent noise present in sensor captures, further affirming the efficacy of our approach in handling challenging real-world scenarios.

### 13 Apr 2024 v2

@@ 4.2.1 Monocular Depth Sequences
To simulate sparse and partial real-world scans, we generated monocular depth sequences by rendering from a fixed camera angle. The size of the point cloud input and the frame length are the same as Sec. 4.1. The qualitative and quantitative comparisons are presented in Fig. 6 and Tab. 2. As seen, our method consistently outperforms all state-of-the-art methods in all metrics and produces more complete surface geometries with more plausible motion tracking. This demonstrates the effectiveness of the motion priors learned by our proposed 4D latent set diffusion in addressing ambiguous data such as partial observations.
@@ 4.2.2 Partial Scan Sequences
To assess the robustness of our method to extremely ambiguous data, we set up a challenging experiment on the D-FAUST [4] dataset. This involved reconstructing whole body motions based on partial point clouds of the upper bodies. This setup creates a highly ambiguous scenario, as the same upper body motion can correspond to many different lower-body. We adopt the same configuration as Sec. 4.1, with a frame length () and input point cloud size (). As the Fig. 7 shows, OFlow [36], LPDC [53], and CaDex [26] face challenges in reconstructing the complete shape, often producing distorted shapes such as broken feet. In contrast, our method excels in reconstructing more complete geometries while achieving temporally coherent tracking. Additionally, our approach present a diverse range of plausible full-body reconstructions that align with the given upper-body scans. The superior performance is primarily attributed to the 4D latent set diffusion. Our diffusion-based method is more capable of tackling the ‘one-to-many’ complexities from extremely partial data.
@@ 4.3 Ablation Study
We conducted ablation studies to validate the effectiveness of each component (See Tab. 3, Fig. 8) under the setting of 4D Shape Completion from monocular noisy depth scans on the D-FAUST [4] dataset.
What is the effect of diffusion model?
4D surface reconstruction from ambiguous observations of noisy, sparse, or partial point clouds is an ill-posed problem. Deterministic models often yield sub-optimal results. We provide comparisons against the variant of one-step regression without diffusion models. As shown in Fig. 8 and Tab. 3, diffusion model uses a probabilistic way to deal with highly ambiguous inputs and generate plausible predictions. Also, diffusion models can handle “one-to-many” problems and generate diverse and creative outputs as shown in Fig. 7.
@@
Instead of using single global latent code, our approach employs 4D latent vector sets. As indicated in Tab. 3, our method significantly outperforms the global latent codes (with ) and captures more accurate 4D motions. It becomes more apparent in unseen identities, demonstrating an enhanced generalization ability.
@@
For the synchronized deformation latent set diffusion, we have integrated the time self-attention layer in our interleaved spatio-temporal attention mechanism. We attempted to remove the layer. However, the results showed a decrease in all metrics, highlighting the effectiveness of the time self-attention layer in maintaining temporal coherence.
@@
To find out the optimal configuration for learning shape and deformation priors within time-varying deformable surfaces, we tried the channel numbers of the shape and deformation latent sets. The experimental results indicated that using channels for 4D latent set diffusion is more suitable, yielding more favorable results.
@@ 5 Conclusion
We present Motion2VecSets, a 4D diffusion model for dynamic surface reconstruction from point cloud sequences. Our method explicitly models the shape and motion distributions of non-rigid objects through an iterative denoising process, using compressed latent sets to generate plausible and diverse outputs. The learned shape and motion diffusion priors can effectively deal with ambiguous observations, including sparse, noisy, and partial data. Compared to encoding shape and deformation with a global latent, our novel 4D latent set representation enables more accurate non-linear motion capture and improves the generalizability to unseen identities and motions. The designed interleaved space and time attention block for synchronized deformation vector sets diffusion enforces temporal-coherent object tracking while reducing computational overhead. Extensive experiments demonstrate our approach’s superiority in reconstructing sparse, partial, and even half-body point clouds on the D-FAUST [4] and DT4D-A [29] datasets, underlining its robustness to various types of imperfect observations. We believe that Motion2VecSets has the potential for future extension into multi-modal domains, such as text-driven 4D generation and RGB video-based 4D reconstruction.
@@ Acknowledgement.
In this supplementary material, we complement our main paper with additional details. It begins with an introduction to the notations used in our paper in Sec. A. The following section, Sec. B, discusses our network architectures, highlighting key design choices and their impact. This is complemented by Sec. C, which provides essential implementation details. Then, we provide more comparisons between our approach and state-of-the-art methods in Sec. D, offering insights into our model’s strengths and improvements. We also provide an additional ablation study to verify the usefulness of latent set diffusion on the setting of 4D shape reconstruction from sparse and noisy point clouds. Finally, in Sec. E, we demonstrate the real-world applicability of our model, underscoring its practical effectiveness.
@@ Appendix A Notations
In this paper, we use the following notations, as summarized in Tab. 4. The symbol represents the sequence length, with the superscript indicating the time step of a frame in the sequence. Our network’s input is a sparse or partial point cloud, denoted as , where each frame comprises points, serving as conditions in reference time. The sets represent shape codes, deformation codes, and conditional codes, respectively, with each -th code denoted by the corresponding lowercase letter. The mesh is symbolized by , where and correspond to the vertices and faces set of the mesh. Furthermore, refers to surface points from ground truth meshes for learning shape, and refers to corresponding sampled points by FPS. We dedicate for representing the sampled surface points for learning deformation, and refer to sampled points by FPS. Moreover, represents the set of query points, where refers to a point inside the point set, and is the output by feeding the query point into the deformation network.
@@
Points from meshes as input for shape autoencoder | |
Points sampled by FPS from meshes as input for shape autoencoder | |
Points from source and target meshes as input for deformation autoencoder | |
Points sampled by FPS from source and target meshes as input for deformation autoencoder | |
@@ B.1 Shape Diffusion
To optimize computational efficiency, we adopt the furthest point sampling (FPS) technique. This method pinpoints crucial points within a point cloud, thereby thinning its density.:
@@
Nosied shape codes are sent to the denoising neural network. The denoiser consists of two blocks: The space-attention block facilitates positional information exchange among codes in different positions, while the condition-attention block injects information from sparse or partial points (conditional input). After repeating this process, we get the denoised latent set .
@@ B.2 Synchronized Deformation Diffusion
As shown in Fig. 10, the deformation diffusion also contains two parts: Deformation Autoencoder Network and Synchronized Deformation Vector Set Diffusion Network.
@@
We employ the cross attention throughout our method, where denotes the query, and denotes the key-value pair. Similarly, we use KL-divergence to retain high-level features to facilitate the learning of the diffusion model. Here we get the deformation latent set :
At inference time, we take the surface points of the source mesh as query points. The interaction between the predicted deformation codes sampled from the Shape Vector Set Diffusion Network, and the positional embedding of query points through cross-attention yields the approximated latent features. Subsequently, a linear layer derives the predicted positions of the target point cloud.
@@
In this stage, the sparse noisy input point clouds pairs green blocks in Fig. 10 are processed with a conditional encoder. The conditional encoder has the same structure with the deformation encoder, through which we get the conditional latent set . Simultaneously, the shape latent set from the deformation encoder is added with Gaussian noise and sent to the denoiser of the diffusion model. As illustrated in the figure, after the space self-attention block, the conditional latent set is injected into the cross-attention block. Following the cross-attention along the temporal domain, and after 18 repetitions, we get the denoised deformation codes.
@@ C.1 Dataset
For shape encoding, our process begins with the application of the Butterfly Subdivision method [67], an interpolating subdivision technique widely used in computer graphics for generating smooth surfaces from polygonal meshes. Following the methodology outlined in [33], we center and rescale all meshes to ensure that the bounding box of each mesh is centered at the origin and the longest edge is normalized to a length of 1. We then uniformly sample 200k points within this normalized cube and compute their occupancy values to determine whether they lie inside or outside the mesh, as detailed in [33]. To enhance our model’s understanding of surface properties, we introduce Gaussian noise at two different levels to the mesh surface points. This process generates 200k near-surface points, whose occupancy values are also computed. These points, combined with an additional 200k points sampled directly from the mesh surface, constitute the input to our network, providing a comprehensive set of data points for both the uniform space and near-surface regions.
@@
Within the context of the partial challenge configuration, a camera was positioned with a fixed viewing angle of 45 degrees directed towards the human or animal subject within the scene. This positioning was undertaken to capture a partial depth observation of the subject, as illustrated in Figure 11. Additionally, in the more restricted partial setting, only half of the human body was observed. This was accomplished by intentionally selecting vertex indices corresponding to the upper body of the SMPL model [31].
@@ Appendix D Additional Results
D.1 Effectiveness of diffusion models
To further demonstrate the efficacy of the diffusion model, we expanded our ablation study to include 4D Shape Reconstruction from sparse and noisy point cloud sequences, utilizing the D-FAUST dataset [4]. This study provides a more comparative analysis against variants of one-step regression integrated with diffusion models, where the input comprises sequences of point clouds of size . Fig. 12 and Tab. 5 present both quantitative and qualitative comparisons. The results clearly indicate that incorporating the diffusion model significantly reduces reconstruction error and yields more precise motion outputs.
D-FAUST | Unseen Motion | Unseen Individual | ||||
IoU↑ | CD↓ | Corr↓ | IoU↑ | CD↓ | Corr↓ | |
W/o. diffusion | 79.1% | 0.070 | 0.076 | 69.3% | 0.089 | 0.098 |
Full | 90.7% | 0.033 | 0.047 | 83.7% | 0.045 | 0.064 |
D.2 4D Shape Reconstruction
In evaluating our model’s performance in 4D shape reconstruction, we adopted a comprehensive set of metrics: Intersection over Union (IoU), chamfer distance, and correspondence distance. These metrics were chosen for their relevance in accurately quantifying shape reconstruction quality. IoU measures the overlap between predicted and ground truth shapes, Chamfer distance quantifies the average closest point distance, and correspondence distance evaluates the accuracy of point-wise correspondences. These metrics align with the standards set in recent studies, such as those by LPDC [53] and OFlow [36]. We present the average metrics across all 17 frames for both D-FAUST [4] and DT4D-A [29] datasets. The quantitative results are shown in Tab. 7, Tab. 8, Tab. 9, and Tab. 10. Our analysis reveals that our model demonstrates superior generalization capabilities in scenarios involving both unseen motions and unseen individuals. Specifically, it outperforms existing approaches, including OFlow [36], LPDC [53], and CaDeX [26], across all evaluated metrics for all time frames. This advancement underscores the efficacy of our approach in handling the complexities of 4D shape reconstruction. Furthermore, qualitative results in Fig. 15 and Fig. 16 show a selection of 8 frames, chosen to represent a diverse range of motions and shapes, from the total 17 to illustrate our model’s performance. In each figure, the upper part displays the results for unseen motion, while the lower part corresponds to unseen individuals. We utilize a chamfer distance error map for visualization, where blue indicates lower error and red signifies higher error. The color-coded error map, computed based on the distance between predicted and ground truth points, provides an intuitive understanding of the model’s accuracy in different scenarios. Our model not only has an overall smaller error on both datasets, but also captures motions more accurately.
[…]
BEHAVE | OFlow | LPDC | CaDeX | Ours |
Chamfer↓ | 0.137 | 0.201 | 0.126 | 0.062 |
D.3 4D Shape Completion
In the more challenging task of 4D shape completion from partial point clouds, our model shows noticable improvements over existing state-of-the-art methods. The quantitative results, as illustrated in Fig. 17 and Fig. 18, demonstrate substantial performance enhancements, with a lower error rate in scenarios involving unseen motions and individuals. This underscores the robustness of our approach. Also, we present more results about the challenging half human setting. The result is demonstrated in the Fig. 14.
Appendix E Real-world Data Test
In this section, we validate our model using data from the real-world BEHAVE dataset [2], employing four Kinect RGB-D cameras to capture RGB color and depth frames. In our case, we utilize a single view from a fixed camera to align with our previous partial scan setting. Similarly, the depth map is back-projected into 3D point cloud, serving as a partial input for our model. In Fig. 13 and Tab. 6, we present a qualitative and quantitative evaluation of our reconstruction process with the corresponding input and reference RGB frame. The results demonstrate the robustness of our model in scenarios characterized by incomplete scans, such as instances where limbs, like the leg or arm, are obscured by structures, such as a grasped object in the hand. This resilience stems from the inherent capabilities of the diffusion model, empowering our model to infer potential structures even in the presence of significant occlusions.
@@ Appendix E Real-world Data Test
Mean | 60.5% | 51.3% | 64.6% | 83.7% | 0.315 | 0.433 | 0.126 | 0.058 | 0.352 | 0.525 | 0.238 | 0.074 |